Helicapter mass 1000kg rises with verticle acceleration of 15ms^2 -passergersweigh 300kg-give magnitude,direction of the total helicapter on surrounding air?

Grade:11

1 Answers

Aman Bansal
12 years ago

Dear Siva,

(a) Force on the floor by the crew and passengers = 300 (15 +10) = 7500 N

(b) Force of the rotor on the surrounding air

=(1000 + 300)(15 + 10) = 32500N downward.

(c) From Newton s third law of motion, the air will have an equal and opposite reaction on the helicopter rotor blades = 32500 N upward.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert